# Positive formula for the product of conjugacy classes on the unitary group

### v1

@@ Abstract.
The convolution product of two generic conjugacy classes of the unitary group $U_{n}$ is described by a probability distribution on the space of central measures which admits a density. Relating the convolution to the quantum Littlewood-Richardson coefficients and using recent results describing those coefficients, we give a manifestly positive formula for this density. In the same flavor as the hive model of Knutson and Tao, this formula is given in terms of a subtraction-free sum of volumes of explicit polytopes. As a consequence, this expression also provides a positive formula for the volume of moduli spaces of $SU_{n}$-valued flat connections on the three-holed two dimensional sphere, which was first given by Witten in terms of an infinite sum of characters.
@@ 1. Introduction
Irreducible characters and conjugacy classes are two dual facets of the space of central measures on a compact group. By the underlying group multiplication, the space of central measure is given a convolution product which allows to expand the product of two characters along the basis of irreducible characters. The same holds in the case of conjugacy classes, for which the expansion is described by a measure on the space of conjugacy classes. It is then a classical fact that this expansion involves in both cases non-negative quantities, for which it is in general hard to give manifestly positive formulas (see [24] and [11] for the two corresponding open problems in the case of the symmetric group $S_{n}$).
@@
In the framework of compact Lie groups, the problem of finding positive combinatorial formulas for the expansion of characters has been solved in the case of the unitary group $U_{n}$ by Littlewood and Richardson [20] and in full generality by Luzstig [21], Kashiwara [13] and Littelmann [19]. The case of conjugacy classes is still open in full generality, and the goal of the present paper is to address this problem in the case of $U_{n}$. Conjugacy classes of $U_{n}$ are indexed by the symmetrized torus $\mathcal{H}=[0,1[^{n}/S_{n}$, and the decomposition of the product of two conjugacy classes $\alpha$ and $\beta$ is defined as a probability distribution $\mathbb{P}[\cdot|\alpha,\beta]$ on $\mathcal{H}$. As for any compact Lie group, $\mathbb{P}[\cdot|\alpha,\beta]$ can be described in a weak sense as a complex weighted sum of characters, each of which being seen as a function on $\mathcal{H}$. In the case of a simple connected Lie group and when the conjugacy classes $\alpha$ and $\beta$ have maximal dimension, in which case $\alpha$ and $\beta$ are called regular, $\mathbb{P}[\cdot|\alpha,\beta]$ has a density $d\mathbb{P}[\cdot|\alpha,\beta]$ with respect to the Lebesgue measure on $\mathcal{H}$. The present paper gives then an explicit positive formula for this density as a subtraction-free sum of volume of some explicit polytopes.
@@
The convolution structure of conjugacy classes is also intimately related to the moduli spaces of $G$-valued flat connections on punctured Riemannian surfaces. This has been first observed by Witten [31], who used this relation to express the volume of such moduli spaces as an infinite sum on characters on the corresponding Lie group $G$. This formula has then been later expressed [23] as an alternating sum of certain volumes which now have an explicit description (see also [28] for an alternative method using residues). An asymptotic formula for this volume in the large $n$ limit has also been given by Levy [18] as a special case of the more evolved study of the Yang-Mills measure. The formula obtained in this paper for the convolution of conjugacy classes directly translates into a manifestly positive expression of the volume of the moduli space of flat connections on the sphere for $G=SU_{n}$ in terms of volumes of explicit polytopes.
@@
Conjugacy classes on $U_{n}$ are naturally related through the exponential map to co-adjoint orbits on the underlying Lie algebra $\mathfrak{u}_{n}$ of Hermitian matrices, which are indexed by the quotient space $\mathbb{R}^{n}/S_{n}$. In the same way as the multiplicative structure yields a convolution product on conjugacy classes of the Lie group $U_{n}$, the additive structure yields a convolution product on co-adjoint orbits of $\mathfrak{u}_{n}$. For co-adjoint orbits of maximal dimension, a positive formula for the density of the convolution product has already been obtained following the work of Knutson and Tao, see [16, 7]. In particular, using those formulas with the results of [23] already allowed to simplify the original expression of Witten for the volume of $SU_{n}$-valued flat connections on the sphere from an infinite sum of complex coefficients to a finite alternate sum of explicit quantities.
@@
The convolution structure on co-adjoint orbits can be seen as a semi-classical limit of the ring of characters on the Lie group [14]. Thanks to this relation, the formula for the positive density in the $\mathfrak{u}_{n}$ case is up to an explicit factor a limit of Littlewood-Richardson coefficients with growing partitions. Such a coefficient had been expressed by Berenstein and Zelevinsky [3] as the number of integers points in a convex polytope whose boundaries depend on the involved partitions. A reformulation of this expression by Knutson and Tao turned this convex polytope into a simple convex body, called discrete hive. Using the discrete hive model, the semiclassical limit yields then a positive formula for the density of convolutions of regular co-adjoint orbits, which is then given by the volume of certain polytopes called continuous hives. Besides their apparent beauty, having positive formulas for such convolution products offered new tools to tackle difficult probabilistic problems concerning invariant measures on the Lie algebras. For example, the hive model for convolution of coadjoint orbits of $U_{n}$ lies at the heart of the recent results of Narayanan and Sheffield [25] on large deviations for the spectrum of sums of conjugation invariant Hermitian random matrices.
@@
When replacing co-adjoint orbits on $\mathfrak{u}_{n}$ by conjugacy classes of $U_{n}$, the convolution product has to be seen instead as a semi-classical limit of the fusion ring of the Lie algebra with growing level [31] (see also [8] for a more probabilistic approach). For each given level, the fusion ring is a specialization of the ring of characters at roots of unity and the structure coefficients of the multiplication of characters in this quotient ring are non-negative integers called fusion coefficients. There exist no general effective combinatorial expression for those coefficients, but in the case of $U_{n}$ this fusion ring is isomorphic to the quantum cohomology ring of Grassmannian on $\mathbb{P}^{1}$ (see [32] for a geometric explanation of this fact). Through this isomorphism, characters of the fusion ring are Schubert classes of the quantum cohomology.
@@
Using a reinterpration of the quantum cohomology of Grassmannian in terms of complex bundles on $\mathbb{P}^{1}$ [27], Buch, Kresch and Tamvakis [6] related the coefficients appearing in the multiplication of Schubert cycles in the quantum cohomology ring of Grassmannian, called quantum Littelwood-Richardson coefficients, to the structure coefficients of the cohomology ring of the two-steps flag variety. Using this relation and proving a conjectural formula of Knutson on the two-steps flag variety, those authors and Purbhoo [5] gave a positive formula for the quantum Littlewood-Richardson coefficients in terms of certain puzzles (see also [4] for an equivariant version). Such puzzles are generalization of puzzles that already appeared in the work of Knutson, Tao and Woodward [17] in a reformulation of the hive model.
@@
In the present paper, we mimick the program achieved in the co-adjoint case to give a positive formula for the density of the convolution of two regular conjugacy classes on $U_{n}$ in terms of volumes of explicit polytopes. One of the main difficulties in the present framework is the absence of convex formulations for the quantum Littlewood-Richardson coefficients. Based on the puzzle expression of [5], we give an expression of these coefficients as the counting of integers points in a finite union of convex bodies indexed by a smaller tiling model. The expression we found for the boundaries of those convex bodies are not fully explicit. Fortunately, they are explicit enough to operate an asymptotic integer counting which simplifies the boundary, yielding then a positive formula for the convolution of two regular conjugacy classes as the volume of a union of convex bodies with explicit boundaries, see Theorem 2.4 and the notations before.
@@
Before this manuscript, several results have been achieved in the description on the convolution of conjugacy classes of $U_{n}$. One of the most important result, independently obtained by Belkale [2] and Agnihotri and Woodward [1] is the description of the support of the convolution of two non necessarily regular conjugacy classes (see also [29]). This support happens to be a convex set with boundary described by certain quantum Littlewood-Richardson coefficient, in the same vein as the solution to the Horn problem given by Klyachko [15] and Knutson and Tao [16]. The regularity of the convolution of two conjugacy classes has also been investigated, with for example results of [33] showing that the convolution of two conjugagy classes has an $L^{2}$ density when the image yields an open set in $\mathcal{H}$.
Finally, let us mention that the exponential maps yields a much deeper relation between the convolution of conjugacy classes and the one of co-adjoint orbits. This has been formalized [9, 30] through the introduction of the wrapping map $\Phi$ from measures on $\mathbb{R}^{n}/S_{n}$ to measures on $\mathcal{H}$. This maps happens to intertwine convolution on both spaces; the wrapping map does not preserve positiveness in general, which prevents from deducing positive formula for the convolution of conjugacy classes from the one for co-adjoint orbits (see also [23] for a concrete deduction of the formula for the conjugacy classes as an alternating sum). However, this map preserves positiveness for measures of $\mathbb{R}^{n}/S_{n}$ having support on $\mathcal{H}$. Up to a scaling, this allows to recover the positive formula for the convolution of co-adjoint orbits as a special case of the formula for conjugacy classes, exactly as Littlewood-Richardson coefficients are particular cases of their quantum version. At the level of volume of moduli spaces of flat connections, this phenomenon had been already noticed by Jeffrey from purely symplectic arguments [12].
@@ Acknowledgement
Both authors are supported by the Agence Nationale de la Recherche funding CORTIPOM ANR-21-CE40-0019. We thank David Garcia-Zelada, Cédric Lecouvey and Thierry Levy for numerous fruitful discussions surrounding this work.
@@ 2. Notations and statement of the result
Fix $n\geq 3$ throughout this manuscript (the case $n=2$ can be handled by direct computation) and denote by $U_{n}$ the unitary group of size $n$. Then, recall that the set of conjugacy classes of $U_{n}$ is homeomorphic to the quotient space $\mathcal{H}=[0,1[^{n}/S_{n}$, where $S_{n}$ acts on $[0,1[^{n}$ by permutation of the coordinates. For $\theta=(\theta_{1}\geq\theta_{2}\geq\dots\geq\theta_{n})\in\mathcal{H}$, denotes by $\mathcal{O}(\theta)$ the corresponding conjugacy class defined by
@@
If the product of two conjugacy classes can be directly defined in the finite group setting, some extra care are needed in the continuous framework and several equivalent definitions are possible depending on the point of view (symplectic, measure theoretic or probabilistic). Let us adopt the probabilistic point of view and say that the product structure on conjugacy classes translates into a convolution product $\ast:\mathcal{M}_{1}(\mathcal{H})\times\mathcal{M}_{1}(\mathcal{H})\rightarrow\mathcal{M}_{1}(\mathcal{H})$ on the space of probability distributions on $\mathcal{H}$ such that for $\theta,\theta^{\prime}\in\mathcal{H}$, $\delta_{\theta}\ast\delta_{\theta^{\prime}}$ is the random variable $p(U_{\theta}U_{\theta^{\prime}})$, where $U_{\theta}$ (resp. $U_{\theta^{\prime}}$) is sampled uniformly on $\mathcal{O}(\theta)$ (resp. $\mathcal{O}(\theta^{\prime})$) and $p:U_{n}\rightarrow\mathcal{H}$ maps an element of $U_{n}$ to the renormalized arguments of its spectrum in $\mathcal{H}$.
Let us denote by $\mathcal{H}_{reg}=\{\theta\in\mathcal{H},\theta_{1}>\theta_{2}>\ldots>\theta_{n}\}$ the set of regular conjugacy classes of $U_{n}$, namely the ones of maximal dimension in $U_{n}$. For $\alpha,\beta\in\mathcal{H}_{reg}$, $\delta_{\alpha}\ast\delta_{\beta}$ admits a density $d\mathbb{P}[\cdot|\alpha,\beta]$ with respect to the Lebesgue measure on $\{\gamma\in\mathcal{H},\sum_{i=1}^{n}\alpha_{i}+\sum_{i=1}^{n}\beta_{i}-\sum_{i=1}^{n}\gamma_{i}\in\mathbb{N}\}$ (see Section 3 for a concrete proof of this classical result).
@@
The main result of the present manuscript is a positive formula for $d\mathbb{P}[\cdot|\alpha,\beta]$ in terms of the volume of polytopes similar to the hive model of Knutson and Tao [16]. Namely, for $\alpha,\beta,\gamma\in\mathcal{H}$ such that $\sum_{i=1}^{n}\alpha_{i}+\sum_{i=1}^{n}\beta_{i}=\sum_{i=1}^{d}\gamma_{i}+d$ for some $d\in\mathbb{N}$, introduce the toric hive
@@
which can be described as a discrete hexagon through the map $(v_{1},v_{2})\mapsto v_{1}+v_{2}e^{i\pi/3}$, see Figure 1 for a particular case and its hexagon representation.
@@
For any set $S$ and any function $f:R_{d,n}\rightarrow S$, we denote by $f^{A}$ (resp $f^{B}$, $f^{C}$) the vector $(f(n+d-i\vee n,i))_{0\leq i\leq n}$ (resp. $(f((d-i)\vee 0,n+d-i\wedge n)_{0\leq i\leq n}$, resp. $(f(n-i,i+d-n\vee 0))_{0\leq i\leq n}$). The vectors $f^{A},f^{B}$ and $f^{C}$ correspond respectively to the east, north-west and south-east boundaries of $R_{d,n}$ through the hexagonale representation, see Figure 2.
@@
Let us call a lozenge of $R_{d,n}$ any sequence $(v^{1},v^{2},v^{3},v^{4})\in R_{d,n}^{4}$ corresponding to one of the three configurations of Figure 3.
@@
A lozenge $(v^{1},v^{2},v^{3},v^{4})$ on which $g$ takes the value $(a,a+1,a+2,a+1)$ for some $a\in\{0,1,2\}$ is called rigid.
@@
Remark that by the boundary condition of a regular labeling, any vertex $v_{4}$ of a rigid lozenge of $g$ can not be on the boundary of $R_{d,n}$, so that the latter is always contained in $Supp(g)$.
@@
As we will see later, for any regular labeling $g$, $Supp(g)$ has cardinal $(n-1)(n-2)/2+3n$ and so is the dimension of $\mathcal{C}_{g}$. As such, we recover the usual dimension of the classical hive cone from [16]. The latter is then a particular case of toric hive cone for $d=0$. An example of a toric hive in the case $n=3,d=1$ is given in Figure 5.
@@
Let $n\geq 3$ and let $\alpha,\beta,\gamma\in\mathcal{H}_{reg}$ be such that $\sum_{i=1}^{n}\alpha_{i}+\sum_{i=1}^{n}\beta_{i}=\sum_{i=1}^{d}\gamma_{i}+d$ with $d\in\mathbb{N}$. Let $g$ be a regular labeling. Then, $P^{g}_{\alpha,\beta,\gamma}$ is the polytope of $\mathbb{R}^{Supp(g)\setminus\partial R_{d,n}}$ consisting of functions in $\mathcal{C}_{g}$ such that
$f^{A}=\left((d-i)^{+}+\sum_{s=1}^{i}\alpha_{s}\right)_{0\leq i\leq n},\,f^{B}=\left(\sum_{s=1}^{n}\alpha_{s}+\sum_{s=1}^{i}\beta_{s}\right)_{0\leq i\leq n},\,f^{C}=\left(d+\sum_{s=1}^{i}\gamma_{s}\right)_{0\leq i\leq n}.$
An example of an element of $P^{g}_{\alpha,\beta,\gamma}$ for $n=3$ and $d=1$ is depicted in Figure 5, for $\alpha=\left(\frac{18}{23}\geq\frac{10}{23}\geq\frac{5}{23}\right)$, $\beta=\left(\frac{13}{23}\geq\frac{6}{23}\geq\frac{2}{23}\right)$ and $\gamma=\left(\frac{20}{23}\geq\frac{9}{23}\geq\frac{2}{23}\right)$.
@@
Our main result gives then a formula for the density of the convolution of regular conjugacy classes in terms of volume of polytopes coming from $\mathcal{C}_{g}$ for regular labeling $g$.
Theorem 2.4.
Let $n\geq 3$ and let $\alpha,\beta,\gamma\in\mathcal{H}_{reg}$ be such that $\sum_{i=1}^{n}\alpha_{i}+\sum_{i=1}^{n}\beta_{i}=\sum_{i=1}^{d}\gamma_{i}+d$ with $d\in\mathbb{N}$. Then,

### v2

@@ Abstract.
The convolution product of two conjugacy classes of the unitary group $U_{n}$ is described by a probability distribution on the space of central measures. Relating this convolution to the quantum cohomology of Grassmannians and using recent results describing the structure constants of the latter, we give a manifestly positive formula for the density of the probability distribution for the product of generic conjugacy classes. In the same flavor as the hive model of Knutson and Tao, this formula is given in terms of a subtraction-free sum of volumes of explicit polytopes. As a consequence, this expression also provides a positive and explicit formula for the volume of $SU_{n}$-valued flat connections on the three-holed two dimensional sphere, which was first given by Witten in terms of an infinite sum of characters.
@@ 1. Introduction
Irreducible characters and conjugacy classes of a compact group are two dual facets of its space of central measures. Thanks to the underlying group multiplication, this space is given a convolution product which allows to expand the product of two characters along the basis of irreducible characters. The same holds in the case of convolution of conjugacy classes, for which the expansion is described by a central measure on the group. It is then a classical fact that both problems involve only non-negative quantities, and it is in general very hard to provide manifestly positive formulas for those expansions (see [29] and [14] for the two corresponding open problems in the case of the symmetric group $S_{n}$).
@@
In the framework of compact Lie groups, the problem of finding positive combinatorial formulas for the expansion of characters has been solved in the case of the unitary group $U_{n}$ by Littlewood and Richardson [24] and in full generality by Luzstig [25], Kashiwara [16] and Littelmann [23]. The case of conjugacy classes is still open in full generality, and the goal of the present paper is to address this problem in the case of $U_{n}$. Conjugacy classes of $U_{n}$ are indexed by the symmetrized torus $\mathcal{H}=[0,1[^{n}/S_{n}$, and the decomposition of the product of two conjugacy classes $\alpha$ and $\beta$ is described by a probability distribution $\mathbb{P}[\cdot|\alpha,\beta]$ on $\mathcal{H}$. As for any compact Lie group, $\mathbb{P}[\cdot|\alpha,\beta]$ can be expressed in a weak sense as a complex weighted sum of characters, each of which being seen as a function on $\mathcal{H}$. When the conjugacy classes $\alpha$ and $\beta$ have maximal dimension, in which case $\alpha$ and $\beta$ are called regular, $\mathbb{P}[\cdot|\alpha,\beta]$ has a density $\mathrm{d}\mathbb{P}[\cdot|\alpha,\beta]$ with respect to the Lebesgue measure on $\mathcal{H}$. The present paper gives then a positive formula for this density as a subtraction-free sum of volumes of some explicit polytopes.
@@
The convolution of conjugacy classes of a compact Lie group $G$ is also intimately related to the moduli spaces of $G$-valued flat connections on punctured Riemannian surfaces. Moduli spaces of flat connections on Riemann surfaces are a central object in algebraic geometry [30] and mathematical physic [2], which appear in particular as the semiclassical limit of two dimensional Yang-Mills measures, [13, 22]. In [36, 37], Witten first proposed a general expression of the volume of such moduli spaces as an infinite sum of characters on the corresponding Lie group $G$. This formula has then been then proved by Jeffrey and Kirwan [15] using localisation techniques from symplectic geometry. The involved infinite series have been later simplified [27] to an alternating sum of volumes of coadjoint orbits, see also [10] for a similar result using the wrapping map. As it has been observed in [37], the study of moduli spaces of flat connections on punctured Riemann surfaces with prescribed holonomies can be reduced to the case of the three punctured sphere, see also [27]. The formula obtained in this paper for the convolution of conjugacy classes directly translates into a simple and manifestly positive expression for the volume of the moduli space of flat connections on the three-punctured sphere for $G=SU_{n}$. This is up to our knowledge the first expression of those volumes as the volume of explicit polytopes.
@@
Let us briefly explain the conceptual path leading to the previous formulas. Conjugacy classes on $U_{n}$ are naturally related through the exponential map to co-adjoint orbits on the underlying Lie algebra $\mathfrak{u}_{n}$ of Hermitian matrices, which are indexed by the quotient space $\mathbb{R}^{n}/S_{n}$. In the same way as the multiplicative structure yields a convolution product on conjugacy classes of the Lie group $U_{n}$, the additive structure yields a convolution product on co-adjoint orbits of $\mathfrak{u}_{n}$. For co-adjoint orbits of maximal dimension, a positive formula for the density of the convolution product has already been obtained following the work of Knutson and Tao, see [19, 8]. Besides their apparent beauty, having positive formulas for such convolution products offered new tools to tackle difficult probabilistic problems concerning invariant measures on the Lie algebras. For example, the positive formulas for the convolution of coadjoint orbits of $U_{n}$ lie at the heart of the recent results of Narayanan and Sheffield [31] on large deviations for the spectrum of sums of conjugation invariant Hermitian random matrices.
@@
The convolution structure on co-adjoint orbits can be seen as a semi-classical limit of the ring of characters of the Lie group [17]. Thanks to this relation, the formula for the positive density in the $\mathfrak{u}_{n}$ case is up to an explicit factor a limit of Littlewood-Richardson coefficients with growing partitions. Such a coefficient had been expressed by Berenstein and Zelevinsky [4] as the number of integers points in a convex polytope whose boundaries depend on the involved partitions. A reformulation of this expression by Knutson and Tao turned this convex polytope into a simple convex body, called discrete hive. Using the discrete hive model, the semiclassical limit yields then a positive formula for the density of convolutions of regular co-adjoint orbits, which the consists of the volume of certain polytopes called continuous hives.
@@
When replacing co-adjoint orbits on $\mathfrak{u}_{n}$ by conjugacy classes of $U_{n}$, the convolution product has to be seen instead as a semi-classical limit of the fusion ring of the Lie algebra with growing level [36] (see also [9] for a more probabilistic approach). For each given level, the fusion ring is a specialization of the ring of characters at roots of unity and the structure coefficients of the multiplication of characters in this quotient ring are non-negative integers called fusion coefficients. There exist no general effective combinatorial expression for those coefficients, but in the case of $U_{n}$ this fusion ring is isomorphic to the quantum cohomology ring of Grassmannians on $\mathbb{P}^{1}$ (see [38] for a geometric explanation of this fact). Through this isomorphism, characters of the fusion ring are Schubert classes of the quantum cohomology.
@@
Using a reinterpretation of the quantum cohomology of Grassmannians in terms of complex bundles on $\mathbb{P}^{1}$ [33], Buch, Kresch and Tamvakis [7] related the coefficients appearing in the multiplication of Schubert classes, called quantum Littelwood-Richardson coefficients, to the structure coefficients of the cohomology ring of the two-step flag variety. Starting from this relation and proving a conjectural formula of Knutson on the two-step flag variety, those authors and Purbhoo [6] gave a positive formula for the quantum Littlewood-Richardson coefficients in terms of certain puzzles (see also [5] for an equivariant version). Such puzzles are generalizations of puzzles that already appeared in the work of Knutson, Tao and Woodward [20] in a reformulation of the hive model. The puzzle approach has since then been systematically used to obtain combinatorial expressions of structure coefficients either in the cohomology or in the K-theory of some partial flag manifolds, see [21].
@@
In this paper, we translate the program achieved in the co-adjoint case to the conjugacy one. One of the main difficulties in the present framework is the absence of convex formulations for the quantum Littlewood-Richardson coefficients. Based on the puzzle formulation of [6], we first express the structure constants of the two-step flag manifold as the counting of integer points in a finite union of convex polytopes indexed by a smaller tiling model, see Theorem 5.3. In order to operate an asymptotic counting, we then turn those convex polytopes into convex bodies, whose boundaries are not fully explicit. Fortunately, the asymptotic integer counting then simplifies the boundaries leading to a formula for the convolution of two regular conjugacy classes as a finite sum of volumes of explicit convex bodies, see Theorem 2.5.
@@
Before this manuscript, several results have been achieved in the description on the convolution of conjugacy classes of $U_{n}$. One of the most important results, independently obtained by Belkale [3] and Agnihotri and Woodward [1], is the description of the support of the convolution of two non necessarily regular conjugacy classes (see also [34]). This support happens to be a convex set with boundaries described by certain quantum Littlewood-Richardson coefficients, in the same vein as the solution to the Horn problem given by Klyachko [18] and Knutson and Tao [19]. The regularity of the convolution of two conjugacy classes has also been investigated, with for example results of [39] showing that the convolution of two conjugagy classes has an $L^{2}$ density when the image is an open set in $\mathcal{H}$.
@@ Acknowledgement
Both authors are supported by the Agence Nationale de la Recherche funding CORTIPOM ANR-21-CE40-0019. We thank David Garcia-Zelada, Cédric Lecouvey and Thierry Lévy for numerous fruitful discussions surrounding this work.
@@ 2. Notations and statement of the result
Fix $n\geq 3$ throughout this manuscript (the case $n=2$ can be handled by direct computation) and denote by $U_{n}$ the unitary group of size $n$. Then, recall that the set of conjugacy classes of $U_{n}$ is homeomorphic to the quotient space $\mathcal{H}=[0,1[^{n}/S_{n}$, where the symmetric group $S_{n}$ acts on $[0,1[^{n}$ by permutation of the coordinates. This quotient space is described by the set of non-increasing sequences of $[0,1[^{n}$. For $\theta=(\theta_{1}\geq\theta_{2}\geq\dots\geq\theta_{n})\in\mathcal{H}$, denote by $\mathcal{O}(\theta)$ the corresponding conjugacy class defined by
@@
The product structure on $U_{n}$ translates into a convolution product $\ast:\mathcal{M}_{1}(\mathcal{H})\times\mathcal{M}_{1}(\mathcal{H})\rightarrow\mathcal{M}_{1}(\mathcal{H})$ on the space of probability distributions on $\mathcal{H}$ such that for $\theta,\theta^{\prime}\in\mathcal{H}$, $\delta_{\theta}\ast\delta_{\theta^{\prime}}$ is the distribution of $p(U_{\theta}U_{\theta^{\prime}})$, where $U_{\theta}$ (resp. $U_{\theta^{\prime}}$) is sampled uniformly on $\mathcal{O}(\theta)$ (resp. $\mathcal{O}(\theta^{\prime})$) and $p:U_{n}\rightarrow\mathcal{H}$ maps an element of $U_{n}$ to its conjugacy class in $\mathcal{H}$.
Let us denote by $\mathcal{H}_{reg}=\{\theta\in\mathcal{H},\theta_{1}>\theta_{2}>\ldots>\theta_{n}\}$ the set of regular conjugacy classes of $U_{n}$, namely the ones of maximal dimension in $U_{n}$. For $\alpha,\beta\in\mathcal{H}_{reg}$, $\delta_{\alpha}\ast\delta_{\beta}$ admits a density $\mathrm{d}\mathbb{P}[\cdot|\alpha,\beta]$ with respect to the Lebesgue measure on $\{\gamma\in\mathcal{H},\sum_{i=1}^{n}\alpha_{i}+\sum_{i=1}^{n}\beta_{i}-\sum_{i=1}^{n}\gamma_{i}\in\mathbb{N}\}$ (see Section 3 for a concrete proof of this classical result).
@@
The main result of the present manuscript is a positive formula for $\mathrm{d}\mathbb{P}[\cdot|\alpha,\beta]$ in terms of the volume of polytopes similar to the hive model of Knutson and Tao [19]. For $0\leq d\leq n$, define the toric hive $R_{d,n}$ as the set
@@
which can be represented as a discrete hexagon through the map $(v_{1},v_{2})\mapsto v_{1}+v_{2}e^{i\pi/3}$, see Figure 1 for a particular case and its hexagonal representation.
@@
For any set $S$ and any function $f:R_{d,n}\rightarrow S$, we denote by $f^{A}$ (resp $f^{B}$, $f^{C}$) the vector $(f((d-i)\vee 0,(n+d-i)\wedge n)_{0\leq i\leq n}$ (resp. $(f(n+d-i\wedge n,i))_{0\leq i\leq n}$, resp. $(f(n-i,i+d-n\vee 0))_{0\leq i\leq n}$). The vectors $f^{A},f^{B}$ and $f^{C}$ correspond respectively to the north-west, east and south-west boundaries of $R_{d,n}$ through the hexagonal representation, see Figure 2.
@@
Let us call a lozenge of $R_{d,n}$ any sequence $(v^{1},v^{2},v^{3},v^{4})\in(R_{d,n})^{4}$ corresponding to one of the three configurations of Figure 3 in the hexagonal representation (in which $|v^{i}-v^{i+1}|=1$ for $1\leq i\leq 3$).
@@
A lozenge $(v^{1},v^{2},v^{3},v^{4})$ for which $(g(v^{1}),g(v^{2}),g(v^{3}),g(v^{4}))=(a,a+1,a+2,a+1)$ for some $a\in\{0,1,2\}$ is called rigid.
@@
By the boundary condition of a regular labeling, any vertex $v_{4}$ of a rigid lozenge of $g$ can not be on the boundary of $R_{d,n}$, so that the latter is always contained in $Supp(g)$.
Remark 2.2.
@@
As we will see later, for any regular labeling $g$, $Supp(g)$ has cardinal $(n-1)(n-2)/2+3n$ and so is the dimension of $\mathcal{C}_{g}$. As such, we recover the usual dimension of the classical hive cone from [19]. The latter is then a particular case of toric hive cone for $d=0$. An example of a toric rhombus concave function in the case $n=3,d=1$ is given in Figure 5.
@@
Let $n\geq 3$ and let $\alpha,\beta,\gamma\in\mathcal{H}_{reg}$ be such that $\sum_{i=1}^{n}\alpha_{i}+\sum_{i=1}^{n}\beta_{i}=\sum_{i=1}^{n}\gamma_{i}+d$ with $d\in\mathbb{N}$. Let $g$ be a regular labeling on $R_{d,n}$. Then, $P^{g}_{\alpha,\beta,\gamma}$ is the polytope of $\mathbb{R}^{Supp(g)\setminus\partial R_{d,n}}$ consisting of functions in $\mathcal{C}_{g}$ such that
$f^{A}=\left(\sum_{s=1}^{n}\beta_{s}+\sum_{s=1}^{i}\alpha_{s}\right)_{0\leq i\leq n},\,f^{B}=\left((d-i)^{+}+\sum_{s=1}^{i}\beta_{s}\right)_{0\leq i\leq n},\,f^{C}=\left(d+\sum_{s=1}^{i}\gamma_{s}\right)_{0\leq i\leq n}.$
An example of an element of $P^{g}_{\alpha,\beta,\gamma}$ for $n=3$ and $d=1$ is depicted in Figure 5, for $\alpha=\left(\frac{13}{23}\geq\frac{6}{23}\geq\frac{2}{23}\right)$, $\beta=\left(\frac{18}{23}\geq\frac{10}{23}\geq\frac{5}{23}\right)$ and $\gamma=\left(\frac{20}{23}\geq\frac{9}{23}\geq\frac{2}{23}\right)$.
@@
Our main result gives then a formula for the density of the convolution of regular conjugacy classes as a sum of volumes of polytopes coming from $\mathcal{C}_{g}$ for regular labeling $g$.
Theorem 2.5 (Probability density for product of conjugacy classes).
Let $n\geq 3$ and let $\alpha,\beta,\gamma\in\mathcal{H}_{reg}$ be such that $\sum_{i=1}^{n}\alpha_{i}+\sum_{i=1}^{n}\beta_{i}=\sum_{i=1}^{n}\gamma_{i}+d$ with $d\in\mathbb{N}$. Then,